If the length of a rectangle is 12 inches. Write an equation to represent the perimeter.
Answer by srinivas.g(540)

let x be the width of rectangle
let y be the perimeter of rectangle
formula : perimeter = 2(length+width)
y = 2(12+x)
y = 2*12+2*x
y = 24+2x
Equation we are looking for y= 24+2x